# Multi-objective Reinforcement Learning with Nonlinear Preferences: Provable Approximation for Maximizing Expected Scalarized Return

### v3

@@ Abstract
1 Introduction
Markov Decision Processes (MDPs) model goal-driven interaction with a stochastic environment, typically aiming to maximize a scalar-valued reward per time-step through a learned policy. Equivalently, this formulation asks the agent to maximize the expected value of a linear function of total reward. This problem can be solved with provable approximation guarantees in polynomial time with respect to the size of the MDP [16, 9, 5, 6, 2].
We extend this framework to optimize a nonlinear function of vector-valued rewards in multi-objective MDPs (MOMDPs), aiming to maximize expected welfare $\mathbb{E}[W(\mathbf{r})]$ where $\mathbf{r}$ is a total reward vector for $d$ objectives. We note that this function is also called the utility or the scalarization function within the multi-objective optimization literature [14]. Unlike deep neural networks used for large state spaces which uses nonlinear function approximators, our nonlinearity lies entirely in the objective function.
@@ 1 Introduction
Nonlinear welfare functions capture richer preferences for agents, such as fairness or risk attitudes. For example, the Nash Social Welfare function $\left(W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}\right)$ reflects a desire for fairness or balance across objectives and diminishing marginal returns [11]. Even single-objective decision theory uses nonlinear utility functions to model risk preferences, like risk aversion with the Von Neumann-Morgenstern utility function [24].
Consider a minimal example where an autonomous taxi robot, Robbie, serving rides in neighborhoods $A$ and $B$, as diagrammed in Figure 1. Each ride yields a reward in its respective dimension.
@@
Figure 1: Taxi Optimization Example
Suppose Robbie starts in neighborhood $A$ and has $t=3$ time intervals remaining before recharging. With no discounting, the Pareto Frontier of maximal (that is, undominated) policies can achieve cumulative reward vectors of $(3,0)$, $(1,1)$, or $(0,2)$. If we want Robbie to prefer the second more balanced or “fair” option of serving one ride in each of the two neighborhoods then Robbie must have nonlinear preferences. However, the second option would maximize the Nash Social Welfare of cumulative reward, for example.
Nonlinear preferences complicate policy computation: Bellman optimality fails, and stationary policies may be suboptimal. Intuitively, a learning agent with fairness oriented preferences to balance objectives should behave differently, even in the same state, depending on which dimension of reward is “worse off.” In the Figure 1 example, one policy to achieve the balanced objective $(1,1)$ is to complete a ride in $A$, then travel from $A$ to $B$, and finally to complete a ride in $B$ – note that this is not stationary with respect to the environment states.
@@
In this work, we ask whether it is possible to describe an approximation algorithm (that is, with provable guarantees to approximate the welfare optimal policy) for MOMDPs with nonlinear preferences that has polynomial dependence on the number of states and actions, as is the case for linear preferences or scalar MDPs. To the best of our knowledge, ours is the first work to give provable guarantees for this problem, compared to other work that focuses exclusively on empirical evaluation of various neural network architectures.
@@ 2 Related Work
Most algorithms for reinforcement learning focus on a single scalar-valued objective and maximizing total expected reward [21]. Classic results on provable approximation and runtime guarantees for reinforcement learning include the E3 algorithm [16]. This result showed that general MDPs could be solved to near-optimality efficiently, meaning in time bounded by a polynomial in the size of the MDP (number of states and actions) and the horizon time. Subsequent results refined the achievable bounds [9, 5, 6]. We extend these results to the multi-objective case with nonlinear preferences.
@@
Multi-objective reinforcement learning optimizes multiple objectives at once. So-called single-policy methods use a scalarization function to reduce the problem to scalar optimization for a single policy, and we follow this line of research. The simplest form is linear scalarization, applying a weighted sum on the Q vector [18, 17, 1, 25, 4].
@@
A more general problem is to optimize the expected value of a potentially nonlinear function of total reward, which may be vector-valued for a multi-objective optimization context. We refer to such a function as a welfare function [7, 13, 20]; it is also commonly referred to as a utility or scalarization function [14, 3]. Other recent work has studied nonlinear objectives, but ours is the first to our knowledge to provide an approximation algorithm with provable guarantees based on a characterization of recursive optimality in this setting. Several others study algorithms with desirable convergenece properties and empirical performance based on conditioning function approximators on accumulated reward but with no approximation guarantees, all of which we see as complementary efforts to our work [20, 13, 10, 19].
@@
[3] describe another model-based algorithm that can compute approximately optimal policies for a general class of monotone and Lipshitz welfare functions, but rather than maximizing the expected welfare, they maximize the welfare of expected rewards (note the two are not equal for nonlinear welfare functions). Other works have formulated fairness in different ways or settings. For example, [15] defines an analogue of envy freeness and [12] studies a per-time-step fairness guarantee. [7] studies welfare maximization in multi-armed bandit problems.
3 Preliminaries
@@
A Multi-objective Markov Decision Process (MOMDP) consists of a finite set $\mathcal{S}$ of states, a starting state $s_{1}\in\mathcal{S}$, [^1] In general we may have a distribution over starting states; we assume a single starting state for ease of exposition. a finite set $\mathcal{A}$ of actions, and probabilities $Pr(s^{\prime}|s,a)\in[0,1]$ that determine the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$. Probabilities are normalized so that $\sum_{s^{\prime}}Pr(s^{\prime}|s,a)=1$.
@@
We also have a vector-valued reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$. Each of the $d$ dimensions of the reward vector corresponds to one of the multiple objectives that are to be maximized. At each time step $t$, the agent observes state $s_{t}\in\mathcal{S}$, takes action $a_{t}\in\mathcal{A}(s_{t})$, and receives a reward vector $\mathbf{R}(s_{t},a_{t})\in\mathbb{R}^{d}$. The environment, in turn, transitions into $s_{t+1}$ with probability $Pr(s_{t+1}|a_{t},s_{t})$.
@@
To make the optimization objective well-posed in MOMDPs with vector-valued rewards, we must specify a scalarization function [14] which we denote as $W:\mathbb{R}^{d}\to\mathbb{R}$. For fair multi-objective reinforcement learning, we think of each of the $d$ dimensions of the reward vector as corresponding to distinct users. The scalarization function can thus be thought of as a welfare function over the users, and the learning agent is a welfare maximizer. Even when $d=1$, a nonlinear function $W$ can be a Von Neumann-Morgenstern utility function [24] that expresses the risk-attitudes of the learning agent, with strictly concave functions expressing risk aversion.
@@
Here we clarify some preliminary assumptions we make about the reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$ and the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$. The restriction to $[0,1]$ is not too restrictive and can be seen simply as a normalization; the substantial assumption is that the rewards are simply finite and bounded. Note from the writing we assume the reward function is deterministic: While it suffices for linear optimization to simply learn the mean of random reward distributions, this does not hold when optimizing the expected value of a nonlinear function. Nevertheless, the environment itself may still be stochastic, as the state transitions may still be random.
@@
This smoothness assumption is necessary for worst-case analysis, but not to run the algorithm.
4 Modeling Optimality
@@
For length $T$ trajectory $\tau$ and discount factor $10\leq\gamma<$ 1, the total discounted reward along $\tau$ is the vector
@@
For ease of exposition we will frequently leave $\gamma$ implicit from context and simply write $\mathbf{R}(\tau)$.
@@
$Pr^{\pi}[\tau]=\pi(a_{1}\mid\tau_{0:0},s_{1})\times\prod_{t=2}^{T}\pi\left(a_{t}~{}\bigg{|}~{}\tau_{1:t-1},s_{t}\right)Pr(s_{t}|s_{t-1},a_{t-1}).$
@@
Note that this objective is not equal to $W\left(\mathbb{E}_{\tau\sim\pi}\bigl{[}\mathbf{R}(\tau)\bigr{]}\right)$, which others have studied [3, 20], for a nonlinear $W$. The former (our objective) is also known as expected scalarized return (ESR) whereas the latter is also known as scalarized expected return (SER) [14]. While SER makes sense for a repeated decision-making problem, it does not optimize for expected welfare for any particular trajectory. For concave $W$, ESR $\leq$ SER by Jensen’s inequality. However, an algorithm for approximating SER does not provide any guarantee for approximating ESR. For example, a policy can be optimal on SER but achieve 0 ESR if it achieves high reward on one or the other of two objectives but never both in the same episode.
@@
We develop an extended form of Bellman optimality capturing this intuition by showing that the optimal value function can be written as a function of current state $s$, accumulated discounted reward $\mathbf{R}(\tau)$, and number of timesteps remaining in the task $t$. The proof is deferred to the appendix.
@@
$V^{*}(s,\mathbf{R}_{acc},t)=\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times V^{*}(s^{\prime},\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a),t-1).$
@@
Note that an approximation algorithm for this discounted finite time horizon problem can also be used as an approximation algorithm for the discounted infinite time horizon problem. Informally, discounting by $\gamma$ with bounded maximum reward implies that the first $T\approx 1/(1-\gamma)$ steps dominate overall returns. We defer the precise formulation of the lower bound on the horizon time and its proof in the appendix.
5 Computing Optimal Policies
Our overall algorithm is Reward-Aware Explore or Exploit (or RAEE for short) to compute an approximately optimal policy, inspired by the classical E3 algorithm [16]. At a high level, the algorithm explores to learn a model of the environment, periodically pausing to recompute an approximately optimal policy on subset of the environment that has been thoroughly explored. We call this optimization subroutine Reward-Aware Value Iteration (or RAVI for short). In both cases, reward-aware refers to the fact that the algorithms compute non-stationary policies in the sense that optimal behavior depends on currently accumulated vector-valued reward, in addition to the current state.
@@ 5 Computing Optimal Policies
A naive algorithm for computing a nonstationary policy would need to consider all possible prior trajectories for each decision point, leading to a runtime complexity containing the term $|\mathcal{S}|^{T}$, exponential in the size of the state space and time horizon. Instead, for a smooth welfare function on a constant number of objectives, our algorithm will avoid any exponential dependence on $|\mathcal{S}|$.
The algorithm is derived from the recursive definition of the optimal multi-objective value function $V^{*}$ in Definition 6, justified by Lemma 1, parameterized by the accumulated discounted reward vector $\mathbf{R_{acc}}$ instead of the prior history. Note that even if rewards were integer, $\mathbf{R_{acc}}$ might not be due to discounting. We must therefore introduce a discretization that maps accumulated reward vectors to points on a lattice, parameterized by some $\alpha\in(0,1)$ where a smaller $\alpha$ leads to a finer discretization but increases the runtime.
@@
For a given discretization precision parameter $\alpha\in(0,1)$, define $f_{\alpha}:\mathbb{R}^{d}\to(\alpha\mathbb{Z})^{d}$ by
@@
In other words, $f$ maps any $d$-dimension vector to its nearest $\alpha$ multiples. We are now prepared to describe the algorithm, which at a high level, uses computes the dynamic program of approximately welfare-optimal value functions conditioned on discretized accumulated reward vectors.
@@
$V(s,\mathbf{R_{acc}},t)\leftarrow\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
@@
$\pi(s,\mathbf{R_{acc}},t)\leftarrow\operatorname*{arg\,max}_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
The asymptotic runtime complexity is $O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha)^{d}\right)$. However, observe that the resulting algorithm is extremely parallelizable: Given solutions to subproblems at $t-1$, all subproblems for $t$ can in principle be computed in parallel. .
The parallel computing implementation of the RAVI algorithm leverages the computational power of GPUs to handle the extensive calculations involved in multi-objective value iteration. Each thread computes the value updates and policy decisions for a specific state and accumulated reward combination, which allows for massive parallelism.
In practice, we observed an empirical speedup of approximately 560 times on a single NVIDIA A100 compared to the CPU implementation on an AMD Ryzen 9 7950x 16-core processor for our experimental settings. This drastic improvement in runtime efficiency makes it feasible to run RAVI on larger and more complex environments, where the computational demands would otherwise be prohibitive.
@@
We begin analyzing the approximation of RAVI with an important structural property: the optimal value function is smooth if the welfare function is smooth. The proof is included in the appendix.
@@
We now present the approximation guarantee of RAVI, that it achieves an additive error that scales with the smoothness of $W$ and the number of time steps remaining. The proof of this lemma is deferred to the appendix.
@@
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in\mathbb{R}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is computed by Algorithm 1.
While we can use any setting of $\alpha$ empirically, this shows that for an approximation guarantee we should set the discretization parameter to $\alpha=\delta_{T\epsilon}/d$ where $\delta_{T\epsilon}$ from the proof of Lemma 3 is sufficient to drive its bound to $\epsilon$, that is, it should be $\alpha=\frac{\delta_{\epsilon}}{T\cdot d}.$
@@
Uniform continuity is a weaker modeling assumption than $L$-Lipshitz continuity for constant $L$. Note that for an $L$-Lipshitz function, the correct value of $\delta_{\epsilon}$ is just $\epsilon/L$, where $\epsilon$ is the desired approximation factor of the algorithm. By taking the gradient, we can see that the $W_{SPF}(\mathbf{x})$ is $d$-Lipshitz for nonnegative $\mathbf{x}$, where $d$ is the dimension of vector reward, so $\delta_{\epsilon}=\epsilon/d$. In practice, as we will see later, this is actually pessimistic: The objective is significantly smoother farther away from the origin, and a discretization that grows increasingly coarse farther from the origin can still efficiently approximate the objective. Nevertheless, Theorem 1 on this example would give us a runtime of
$O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha_{T\epsilon})^{d}\right)=O\left(|\mathcal{S}||\mathcal{A}|\left(\frac{T^{2}\cdot d^{2}}{\epsilon}\right)^{d}\right).$
To further simplify, if one takes the number of actions $|\mathcal{A}|$, discount factor $\gamma$, and dimension $d$ to be small constants, then the asymptotic dependence of the runtime (in our running example) is

### v4

@@ Abstract.
Key words and phrases:
Multi-objective Reinforcement Learning, Nonlinear Optimization, Algorithmic Fairness, Approximation Algorithms
11footnotetext: Equal contribution.
1. Introduction
Markov Decision Processes (MDPs) model goal-driven interaction with a stochastic environment, typically aiming to maximize a scalar-valued reward per time-step through a learned policy. Equivalently, this formulation asks the agent to maximize the expected value of a linear function of total reward. This problem can be solved with provable approximation guarantees in polynomial time with respect to the size of the MDP Kearns and Singh (2002); Brafman and Tennenholtz (2003); Auer et al. (2008); Azar et al. (2017); Agarwal et al. (2021).
We extend this framework to optimize a nonlinear function of vector-valued rewards in multi-objective MDPs (MOMDPs), aiming to maximize expected welfare $\mathbb{E}[W(\mathbf{r})]$ where $\mathbf{r}$ is a total reward vector for $d$ objectives. We note that this function is also called the utility or the scalarization function within the multi-objective optimization literature Hayes et al. (2023). Unlike the deep neural networks commonly used as function approximators for large state spaces, our nonlinearity lies entirely in the objective function.
@@ 1. Introduction
Nonlinear welfare functions capture richer preferences for agents, such as fairness or risk attitudes. For example, the Nash Social Welfare function $\left(W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}\right)$ reflects a desire for fairness or balance across objectives and diminishing marginal returns Caragiannis et al. (2019). Even single-objective decision theory uses nonlinear utility functions to model risk preferences, like risk aversion with the Von Neumann-Morgenstern utility function Von Neumann and Morgenstern (1947).
Consider a minimal example with an autonomous taxi robot, Robbie, serving rides in neighborhoods $A$ and $B$, as diagrammed in Figure 1. Each ride yields a reward in its respective dimension.
@@
Figure 1. Taxi Optimization Example
Suppose Robbie starts in neighborhood $A$ and has $t=3$ time intervals remaining before recharging. With no discounting, the Pareto Frontier of maximal (that is, undominated) policies can achieve cumulative reward vectors of $(3,0)$, $(1,1)$, or $(0,2)$ by serving $A$ alone, $A$ and $B$, or $B$ alone respectively. If we want Robbie to prefer the second more balanced or “fair” option of serving one ride in each of the two neighborhoods then Robbie must have nonlinear preferences. That is, for any choice of weights on the first and second objective, the simple weighted average would prefer outcomes (3, 0) or (0, 2). However, the second option would maximize the Nash Social Welfare of cumulative reward, for example.
Nonlinear preferences complicate policy computation: Bellman optimality fails, and stationary policies may be suboptimal. Intuitively, a learning agent with fairness-oriented preferences to balance objectives should behave differently, even in the same state, depending on which dimension of reward is “worse off.” In the Figure 1 example, one policy to achieve the balanced objective $(1,1)$ is to complete a ride in $A$, then travel from $A$ to $B$, and finally to complete a ride in $B$ – note that this is not stationary with respect to the environment states.
@@
In this work, we ask whether it is possible to describe an approximation algorithm (that is, with provable guarantees to approximate the welfare optimal policy) for MOMDPs with nonlinear preferences that has polynomial dependence on the number of states and actions, as is the case for linear preferences or scalar MDPs. To the best of our knowledge, ours is the first work to give provable guarantees for this problem, compared to other work that focuses on empirical evaluation of various neural network architectures.
@@ 2. Related Work
Most reinforcement learning algorithms focus on a single scalar-valued objective and maximizing total expected reward Sutton and Barto (2018). Classic results on provable approximation and runtime guarantees for reinforcement learning include the E3 algorithm Kearns and Singh (2002). This result showed that general MDPs could be solved to near-optimality efficiently, meaning in time bounded by a polynomial in the size of the MDP (number of states and actions) and the horizon time. Subsequent results refined the achievable bounds Brafman and Tennenholtz (2003); Auer et al. (2008); Azar et al. (2017). We extend these results to the multi-objective case with nonlinear preferences.
@@
Multi-objective reinforcement learning optimizes multiple objectives at once. So-called single-policy methods use a scalarization function to reduce the problem to scalar optimization for a single policy, and we follow this line of research. The simplest form is linear scalarization, applying a weighted sum to the Q vector Moffaert et al. (2013); Liu et al. (2014); Abels et al. (2019); Yang et al. (2019); Alegre et al. (2023).
@@
A more general problem is to optimize the expected value of a potentially nonlinear function of the total reward, which may be vector-valued in a multi-objective optimization context. We refer to such a function as a welfare function Barman et al. (2023); Fan et al. (2023); Siddique et al. (2020), which is also commonly referred to as a utility or scalarization function Hayes et al. (2023); Agarwal et al. (2022). Recent works have explored nonlinear objectives; however, to our knowledge, ours is the first to provide an approximation algorithm with provable guarantees (on the approximation factor), for the expected welfare, leveraging a characterization of recursive optimality in this setting. Several other studies focus on algorithms that demonstrate desirable convergence properties and strong empirical performance by conditioning function approximators on accumulated reward, but without offering approximation guarantees Siddique et al. (2020); Fan et al. (2023); Cai et al. (2023); Reymond et al. (2023). Complementary to our approach, Reymond et al. (2022) uses Pareto Conditioned Networks to learn policies for Pareto-optimal solutions by conditioning policies on a preference vector. Lin et al. (2024) presents an offline adaptation framework that employs demonstrations to implicitly learn preferences and safety constraints, aligning policies with inferred preferences rather than providing theoretical guarantees.
@@
Agarwal et al. (2022) describe another model-based algorithm that can compute approximately optimal policies for a general class of monotone and Lipschitz welfare functions, but rather than maximizing the expected welfare, they maximize the welfare of expected rewards (note the two are not equal for nonlinear welfare functions). Other works have formulated fairness in different ways or settings. For example, Jabbari et al. (2017) defines an analogue of envy freeness and Deng et al. (2023) studies a per-time-step fairness guarantee. Barman et al. (2023) studies welfare maximization in multi-armed bandit problems. Röpke et al. (2023) explores the concept of distributional multi-objective decision making for managing uncertainty in multi-objective environments.
Risk-sensitive RL approaches address scalar objectives by incorporating risk measures to minimize regret or control reward variance over accumulated rewards Bäuerle and Ott (2011); Bellemare et al. (2023); Bastani et al. (2022). While these approaches offer valuable tools for managing reward variability, their guarantees are primarily in terms of regret minimization or achieving bounded variance. In contrast, our work provides stronger theoretical assurances in terms of the approximation ratio on the expected welfare for multi-objective optimization. This difference in focus underscores the robustness of our method, which provides guarantees that extend beyond the risk-sensitive regime to cover complex, multi-dimensional utility functions Fan et al. (2023); Brafman and Tennenholtz (2003); Azar et al. (2017). Such problems remain computationally significant even in a deterministic environment where the notion of risk may not apply.
Lastly, while our single-agent setup with multiple objectives shares some aspects with multi-agent reinforcement learning (MARL), the objectives differ significantly. Much of the MARL literature has focused on cooperative reward settings, often using value-decomposition techniques like VDN and QMIX Sunehag et al. (2017); Rashid et al. (2018) or actor-critic frameworks to align agent objectives under a centralized training and decentralized execution paradigm Foerster et al. (2018). In contrast, our work parallels the more general Markov game setting, where each agent has a unique reward function and studies nonlinear objectives that require computational methods beyond linear welfare functions often assumed in MARL. While MARL research frequently uses linear summations of agent rewards, we demonstrate approximation guarantees for optimizing general, non-linear functions of multiple reward vectors, a distinct contribution in a single-agent setting Jabbari et al. (2017); Agarwal et al. (2022); Von Neumann and Morgenstern (1947).
3. Preliminaries
@@
A finite Multi-objective Markov Decision Process (MOMDP) consists of a finite set $\mathcal{S}$ of states, a starting state $s_{1}\in\mathcal{S}$,***In general we may have a distribution over starting states; we assume a single starting state for ease of exposition. a finite set $\mathcal{A}$ of actions, and probabilities $Pr(s^{\prime}|s,a)\in[0,1]$ that determine the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$. Probabilities are normalized so that $\sum_{s^{\prime}}Pr(s^{\prime}|s,a)=1$.
@@
We have a finite vector-valued reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$. Each of the $d$ dimensions of the reward vector corresponds to one of the multiple objectives that are to be maximized. At each time step $t$, the agent observes state $s_{t}\in\mathcal{S}$, takes action $a_{t}\in\mathcal{A}(s_{t})$, and receives a reward vector $\mathbf{R}(s_{t},a_{t})\in[0,1]^{d}$. The environment, in turn, transitions to $s_{t+1}$ with probability $Pr(s_{t+1}|a_{t},s_{t})$.
@@
To make the optimization objective well-posed in MOMDPs with vector-valued rewards, we must specify a scalarization function Hayes et al. (2023) which we denote as $W:\mathbb{R}^{d}\to\mathbb{R}$. For fair multi-objective reinforcement learning, we think of each of the $d$ dimensions of the reward vector as corresponding to distinct users. The scalarization function can thus be thought of as a welfare function over the users, and the learning agent is a welfare maximizer. Even when $d=1$, a nonlinear function $W$ can be a Von Neumann-Morgenstern utility function Von Neumann and Morgenstern (1947) that expresses the risk-attitudes of the learning agent, with strictly concave functions expressing risk aversion.
@@
Here we clarify some preliminary assumptions we make about the reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$ and the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$. The restriction to $[0,1]$ is simply a normalization for ease of exposition; the substantial assumption is that the rewards are finite and bounded. Note from the writing we assume the reward function is deterministic: While it suffices for linear optimization to simply learn the mean of random reward distributions, this does not hold when optimizing the expected value of a nonlinear function. Nevertheless, the environment itself may still be stochastic, as the state transitions may still be random.
@@
The smoothness assumption seems necessary to give worst-case approximation guarantees as otherwise arbitrarily small changes in accumulated reward could have arbitrarily large differences in welfare. However, we note that this is still significantly more general than linear scalarization which is implicitly smooth. Practically speaking, our algorithms can be run regardless of the assumed level of smoothness; a particular smoothness is necessary just for the worst-case analysis.
4. Modeling Optimality
@@
For length $T$ trajectory $\tau$ and discount factor $0\leq\gamma\leq 1$, the total discounted reward along $\tau$ is the vector
@@
For ease of exposition we will frequently leave $\gamma$ implicit from context and simply write $\mathbf{R}(\tau)$. $\gamma<1$ is necessary for the infinite horizon setting. In the experiments with a finite-horizon task we use $\gamma=1$ for simplicity.
@@
$Pr^{\pi}[\tau]=\pi(a_{1}\mid\tau_{0:0},s_{1})\times\prod_{t=2}^{T}\pi\left(a_{t}\leavevmode\nobreak\ \bigg{|}\leavevmode\nobreak\ \tau_{1:t-1},s_{t}\right)Pr(s_{t}|s_{t-1},a_{t-1}).$
@@
Note that this objective is not equal to $W\left(\mathbb{E}_{\tau\sim\pi}\bigl{[}\mathbf{R}(\tau)\bigr{]}\right)$, which others have studied Agarwal et al. (2022); Siddique et al. (2020), for a nonlinear $W$. The former (our objective) is also known as expected scalarized return (ESR) whereas the latter is also known as scalarized expected return (SER) Hayes et al. (2023). While SER makes sense for a repeated decision-making problem, it does not optimize for expected welfare for any particular trajectory. For concave $W$, ESR $\leq$ SER by Jensen’s inequality. However, an algorithm for approximating SER does not provide any guarantee for approximating ESR. For example, a policy can be optimal on SER but achieve 0 ESR if it achieves high reward on one or the other of two objectives but never both in the same episode.
@@
We develop an extended form of Bellman optimality capturing this intuition by showing that the optimal value function can be written as a function of current state $s$, accumulated discounted reward $\mathbf{R}(\tau)$, and number of timesteps remaining in the task $t$. The proof is included in the appendix. At a high level as a sketch, the argument proceeds inductively on $t$ where the base case follows by definition and the inductive step hinges on showing that the the expectation over future trajectories can be decomposed into an expectation over successor states and subsequent trajectories despite the nonlinear $W$.
@@
$V^{*}(s,\mathbf{R}_{acc},t)=\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\cdot V^{*}(s^{\prime},\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a),t-1).$
@@
Note that an approximation algorithm for this discounted finite time horizon problem can also be used as an approximation algorithm for the discounted infinite time horizon problem. Informally, discounting by $\gamma$ with bounded maximum reward implies that the first $T\approx 1/(1-\gamma)$ steps dominate overall returns. We defer the precise formulation of the lower bound on the horizon time and its proof in the appendix (Lemma 13).
Necessity of Conditioning on Remaining Timesteps $t$
We illustrate the necessity of conditioning the optimal value function on the remaining timesteps $t$ using a counterexample in the appendix.
5. Computing Optimal Policies
Our overall algorithm is Reward-Aware Explore or Exploit (or RAEE for short) to compute an approximately optimal policy, inspired by the classical E3 algorithm Kearns and Singh (2002). At a high level, the algorithm explores to learn a model of the environment, periodically pausing to recompute an approximately optimal policy on subset of the environment that has been thoroughly explored. We call this optimization subroutine Reward-Aware Value Iteration (or RAVI for short). In both cases, reward-aware refers to the fact that the algorithms compute non-stationary policies in the sense that optimal behavior depends on currently accumulated vector-valued reward, in addition to the current state.
@@ 5. Computing Optimal Policies
A naive algorithm for computing a non-stationary policy would need to consider all possible prior trajectories for each decision point, leading to a runtime complexity containing the term $|\mathcal{S}|^{T}$, exponential in the size of the state space and time horizon. Instead, for a smooth welfare function on a constant number of objectives, our algorithm will avoid any exponential dependence on $|\mathcal{S}|$.
The algorithm is derived from the recursive definition of the optimal multi-objective value function $V^{*}$ in Definition 7, justified by Lemma 6, parameterized by the accumulated discounted reward vector $\mathbf{R_{acc}}$ instead of the prior history. Note that even if the rewards were integers, $\mathbf{R_{acc}}$ might not be due to discounting. We must therefore introduce a discretization that maps accumulated reward vectors to points on a lattice, parameterized by some $\alpha\in(0,1)$ where a smaller $\alpha$ leads to a finer discretization but increases the runtime.
@@
For a given discretization precision parameter $\alpha\in\mathbb{R}^{+}$, define $f_{\alpha}:\mathbb{R}^{d}\to(\alpha\mathbb{Z})^{d}$ by
@@
In other words, $f_{\alpha}$ maps any $d$-dimensional vector to the largest vector in $(\alpha\mathbb{Z})^{d}$ that is less than or equal to the input vector, effectively rounding each component down to the nearest multiple of $\alpha$. We now describe the algorithm, which at a high level computes the dynamic program of approximately welfare-optimal value functions conditioned on discretized accumulated reward vectors.
Remark. Since we model the per-step reward as normalized to at most 1, we describe $\alpha$ as lying within $(0,1)$. However, the algorithm itself is well-defined for larger values of alpha (coarser than the per-step reward) and during implementation and experiments, we consider such larger $\alpha$, beyond what might give worst-case guarantees but still observing strong empirical performance.
@@
$V\left(s,\mathbf{R_{acc}},t\right)\leftarrow\max_{a}\sum_{s^{\prime}}Pr\left(s^{\prime}|s,a\right)V\left(s^{\prime},\varphi(s,\mathbf{R}_{acc},a),t-1\right)$
@@
$\pi\left(s,\mathbf{R_{acc}},t\right)\leftarrow\operatorname*{arg\,max}_{a}\sum_{s^{\prime}}Pr\left(s^{\prime}|s,a\right)V\left(s^{\prime},\varphi(s,\mathbf{R}_{acc},a),t-1\right)$
9:              where $\varphi(s,\mathbf{R}_{acc},a):=f_{\alpha}\left(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)\right)$
The asymptotic runtime complexity is $O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha)^{d}\right)$. However, observe that the resulting algorithm is extremely parallelizable: Given solutions to subproblems at $t-1$, all subproblems for $t$ can in principle be computed in parallel. A parallel implementation of RAVI can leverage GPU compute to handle the extensive calculations involved in multi-objective value iteration. Each thread computes the value updates and policy decisions for a specific state and accumulated reward combination, which allows for massive parallelism. In practice, we observed an empirical speedup of approximately 560 times on a single NVIDIA A100 compared to the CPU implementation on an AMD Ryzen 9 7950x 16-core processor for our experimental settings. This drastic improvement in runtime efficiency makes it feasible to run RAVI on larger and more complex environments, where the computational demands would otherwise be prohibitive.
@@
We begin analyzing the approximation of RAVI by showing an important structural property: the optimal value function will be smooth as long as the welfare function is smooth. The proof is included in the appendix.
@@
We now present the approximation guarantee of RAVI, that it achieves an additive error that scales with the smoothness of $W$ and the number of remaining time steps. The proof of this lemma is deferred to the appendix.
@@
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in\mathbb{R}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is computed by Algorithm 1 using $\alpha_{\epsilon}$.
While we can use any setting of $\alpha$ empirically, this shows that for an approximation guarantee we should set the discretization parameter to $\alpha=\delta_{T\epsilon}/d$ where $\delta_{T\epsilon}$ from the smoothness of the welfare function is sufficient to drive its bound to $\epsilon$, that is, it should be $\alpha=\frac{\delta_{\epsilon}}{T\cdot d}.$
@@
By taking the gradient, we can see that $W_{\text{SPF}}(\mathbf{x})$ is $d$-Lipschitz on $\mathbf{x}\geq 0$, so we may pick $\delta_{\epsilon}=\epsilon/d$ to satisfy the uniform continuity requirement in Theorem 11. Uniform continuity is a weaker modeling assumption than $L$-Lipshitz continuity for constant $L$. Note that for an $L$-Lipshitz function, the correct value of $\delta_{\epsilon}$ is just $\epsilon/L$, where $\epsilon$ is the desired approximation factor of the algorithm.
Plugging $\delta_{\epsilon}$ into the runtime and recalling that $\alpha_{T\epsilon}=\tfrac{\delta_{T\epsilon}}{d}$ and $\delta_{T\epsilon}=\delta_{\epsilon}/T$ up to constant factors, we get
$O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha_{T\epsilon})^{d}\right)=O\left(|\mathcal{S}|^{2}|\mathcal{A}|\left(\frac{T^{2}\cdot d^{2}}{\epsilon}\right)^{d}\right).$
To further simplify, if one takes the number of actions $|\mathcal{A}|$, discount factor $\gamma$, and dimension $d$ to be constants, then the asymptotic dependence of the runtime (in our running example) is